OpenStudy (anonymous):

how can i solve these with substitution? x+y=12 2x+5y=27

OpenStudy (anonymous):

solve for y in the first equation (y = -x+12) then plug into second equation: 2x + 5(-x+12) = 27

OpenStudy (anonymous):

isolate variable onto one side and constants on the other..we get -3x = -33 so x = 11

OpenStudy (anonymous):

@SamanthaScarecrow have you got it or i'll show you how to solve..??

OpenStudy (anonymous):

2x + 5(-x + 12) = 27 2x - 5x + 60 = 27 -3x + 60 = 27 -3x = -33 x = 11

OpenStudy (anonymous):

i have 2x+5(-x+12)=27 then i distributed so i got 2x+5x+60=27 , but because there was a minus sign infront of the x in the parenthesis i don't know if i should make it 2x-5x+60=27 or if it should be 2+5x+60=27.. so if you could show me how to do it from there @Kashan i would be very grateful c:

OpenStudy (anonymous):

check ur signs when distributing!

OpenStudy (anonymous):

look at my work above :)

OpenStudy (anonymous):

2x-5x+60=27

OpenStudy (anonymous):

ohhhh! @sportegirl i see how you did it there! okay i keep getting messed up with those negatives >< but thank you to both of you for helping me! <3

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

2x-5x= -3x so, -3x+60= 27 -3x = 27-60 -3x = -33 x= -33/ -3 x= 11

OpenStudy (anonymous):

got it now @SamanthaScarecrow ??

OpenStudy (anonymous):

i do c: thank you

OpenStudy (anonymous):

no problem.. :)